Louisiana State University Health Sciences Center Administration & Finance
 

Friday, September 08, 2006

  • Dental School:
    • Pipe insulation labeling/identification continues. The contractors have completed the labeling/identification of the steam condensate pipe and are continuing to label the remaining water pipes in the basement.
    • Soft Water System piping continues as the contractors are replacing the soft water piping in the Power Plant. All three of the soft water tanks have had the pipes & gauges changed. The contractor has filled the tanks & has added the salt mixture. All that remains is the start-up & testing of the system.
    • Mold Remediation of heavily molded rooms (6413B, 6409, & 7409A) is complete. The contractors have installed the new drywall and have completed the application of new wall finishes for these areas (i.e. paint and cove base). The contractors have removed the safety barriers around these areas & are awaiting inspection.
    • The work remaining on the Emergency Generator is the swapping of connections from the temporary-temporary to the temporary-permanent.
    • The Elevator contractor is still awaiting additional parts & new elevator doors prior to resuming work on the elevators. Additional parts are expected to arrive within a couple of weeks.
    • The Medical Gas System is being rescheduled for testing pending the arrival of gas cylinders which are to be provided by LSUHSC. LSUHSC is working to procure the cylinders through the Purchasing Department.
    • The contractor pressure washed the slab & cleared the floor drain strainers in the mechanical room of the basement.
  • 1542 Tulane Ave:
    • Wink Engineering is working on plans to provide power and replace mechanical components that will make the building more suitable for occupancy.
    • The contractor is in the process of changing out the motors and the controllers for the fire pumps in the basement.
    • Ordes Electric is also labeling the new panel boxes that they have finished connecting to the transformer and the equipment they will feed.
    • Hi-pot testing of the cables for the 4th floor riser continues.
  • Utility Plant:
    • We are still waiting for the Variable Speed Drive Motor to be delivered.
    • The repair of the pump for the chillers has been delayed by the subcontractors and is a week or more out from being ready for installation.
  • Campus Window Replacement:
    • The additional windows are still several weeks out
    • The glass doors for Residence Hall were delivered and they will be installed the first part of next week
    • The six windows at 1542 have been replaced